# The infinitesimal structure of manifolds with non-continuous Riemannian metrics

### v1

@@ Abstract
This paper investigates the failure of certain metric measure spaces to be infinitesimally Hilbertian or quasi-Riemannian manifolds, by constructing examples arising from a manifold $M$ endowed with a Riemannian metric $g$ that is possibly discontinuous, with $g,g^{-1}\in L^{\infty}_{\mathrm{loc}}$ and $g\in W^{1,p}_{\mathrm{loc}}$ for $p<\dim M-1$. footnotetext: MSC2020: 53C20, 53C23
@@
Acknowledgements. The author gratefully acknowledges Andrea Mondino for his supervision and guidance, and thanks Alessandro Cucinotta for valuable discussions. The author is supported by a postgraduate scholarship from the Mathematical Institute at the University of Oxford. Further financial support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813, which funded the 2025 Trimester Program “Metric Analysis” in Bonn, is also acknowledged with gratitude.
@@ 1 Introduction
Manifolds with (semi)-Riemannian metrics of low regularity (in particular below $C^{2}$) have been of growing interest and importance in the field of geometric analysis. They may arise from gluing constructions, or in the case of Lorentzian signature (,+,,+)(-,+$\ldots$,+)( - , + , … , + ), serve to understand physically relevant phenomena from general relativity (see for instance [5]).
One approach to generalise Riemannian manifolds is to drop any underlying smoothness assumption and study metric measure spaces. A metric measure space is a triplet $(X,\mathsf{d},\mathfrak{m})$, where $(X,\mathsf{d})$ is a complete and separable metric space and $\mathfrak{m}$ is a $\sigma$-finite Borel measure on $X$ (playing the role of the volume measure). For a similar generalisation of Lorentzian manifolds see [22, 6].
@@
In this paper however, we will focus on the generalised first order calculus on metric measure spaces. It emerged through the search for a criterion on metric measure spaces that distinguishes those arising from Finslerian structures from metric measure spaces arising from Riemannian structures. This lead to the study of Sobolev calculus on metric measure spaces. After the approaches from Cheeger [7], Hajłasz [19], and Shanmungalingam [32], Ambrosio, Gigli and Savaré introduced a Sobolev calculus based on the notion of test plans in [2] that they proved to be equivalent to the approaches in [7] and [32], leading to the definition of the Sobolev space $W^{1,2}(X)$, and a Laplace operator.
In [14], Gigli studied conditions under which the Laplacian is linear and proposed the notion of infinitesimal Hilbertianity. A metric measure space is infinitesimally Hilbertian, if the associated Sobolev space $W^{1,2}(X)$ is a Hilbert space, or equivalently if the Cheeger energy (a generalisation of the Dirichlet energy to metric measure spaces) is a quadratic form. Then the Laplacian is automatically linear and the theory on Dirichlet forms provides various tools to better understand the Laplacian and the heat flow.
Examples of infinitesimally Hilbertian metric measure spaces include geodesically complete weighted Riemannian manifolds [26], locally $\mathrm{CAT}(\kappa)$-spaces equipped with a positive and locally finite Radon measure [11], and sub-Riemannian manifolds equipped with an arbitrary Radon measure [23]. Relevant notions to study the Cheeger energy, are the metric speed of absolutely continuous curves ([1], [21]) and the slope of Lipschitz continuous functions. In [3], Ambrosio, Gigli, and Savaré proposed the Riemannian curvature dimension condition, $\mathsf{RCD}(K,\infty)$, a refinement of the curvature dimension condition which combines the $\mathsf{CD}(K,\infty)$-condition with infinitesimal Hilbertianity. The $\mathsf{RCD}(K,N)$-condition for $N<\infty$ was proposed by Gigli in [13]. Using this terminology, an earlier result obtained by Ohta and Sturm in [30] states that $(\mathbb{R}^{n},\norm{\cdot},\mathcal{L}^{n})$ is an $\mathsf{RCD}(0,N)$-space if and only if $\norm{\cdot}$ arises from an inner product, hence in that case infinitesimal Hilbertianity indeed rules out Riemannian structures among Finslerian ones.
@@
However, there are several relevant classes of Riemannian metrics that are non-continuous, for instance the Geroch-Traschen class of metrics [12], i.e. those Riemannian metrics $g$ such that $g,g^{-1}\in L^{\infty}_{\rm loc}(M)$ and $g\in W^{1,2}_{\rm loc}(M)$. These regularity assumptions are minimal to compute the distributional Ricci curvature on $M$ (see [18], [16]) and have been studied in both Riemannian (see for example [17]) and Lorentzian signature (see [33]). Moreover, in the context of potential theory for uniformly elliptic operators, the even broader class of (Riemannian) metrics $g$ such that $g,g^{-1}\in L^{\infty}_{\rm loc}(M)$ has been studied in [31], [34], [29] and [9].
If the metric tensor is only measurable, Norris (cf. [29]) and De Cecco-Palmieri (cf. [9]) defined a distance $\mathsf{d}_{g}$ that avoids the difficulty that piecewise smooth curves $\gamma$ might lie on a null set where $g$ is not well-defined (see Section 3).
@@
(Under which regularity assumptions on $g$) does $(M,\mathsf{d}_{g},g)$ satisfy that for any Lipschitz continuous function $f:M\to\mathbb{R}$ its slope equals the norm of its gradient almost everywhere, i.e. $|Df|=\limsup_{y\to x}\frac{|f(y)-f(x)|}{\mathsf{d}_{g}(x,y)}=|\nabla_{g}f|_{g}$ almost everywhere?
@@
(Under which regularity assumptions on $g$) does the metric speed of “almost every” absolutely continuous curves equal the norm of its derivative, i.e. $\lim_{h\to 0}\frac{\mathsf{d}_{g}(\gamma_{t},\gamma_{t+h})}{|h|}=|\dot{\gamma}_{t}|_{g}$ for almost every $t\in[0,1]$?
@@
(Under which regularity assumptions on $g$) is $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ infinitesimally Hilbertian?
@@
It is already known that assuming $g\in C^{0}$ is sufficient for the second and third question to have positive answers (see [4], [28]). Moreover, it is known that for $g\in C^{0}$ and $f\in C^{1}$, the slope $|Df|$ of $f$ equals the norm of its gradient [28]. In this paper, we will provide examples showing the necessity of those regularity assumptions.
Outline of the paper and main results. In Section 2, we will recall some basic notions on the first order calculus on metric measure spaces and their explicit descriptions in the case of a manifold equipped with a continuous Riemannian metric. In Section 3, we recall the definition of the distance in the case of a non-continuous Riemannian metric by Norris and De Cecco-Palmieri, and briefly study absolutely continuous curves and test plans in that setting. In Section 4, we will construct Riemannian metrics $g$ with $g,g^{-1}\in L^{\infty}_{\rm loc}$ such that $(M,\mathsf{d}_{g},g)$ is not a quasi-Riemannian manifold and such that $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is not infinitesimally Hilbertian. We start with metrics that only satisfy $g,g^{-1}\in L^{\infty}_{\rm loc}$, providing a first example of a non-quasi Riemannian manifold in Subsection 4.1 and a first example of a non-infinitesimally Hilbertian space in Subsection 4.2. In Subsection 4.3, we refine the previous examples to a Sobolev metric $g$ such that $g,g^{-1}\in W^{1,p}_{\rm loc}\cap L^{\infty}_{\rm loc}$ for some $p<\dim M-1$. Questions 1 and 2 are then addressed in Subsection 4.3.1, where we prove:
@@
For $d\geq 3$, $p\in[1,d-1)$ there exists a $d$-dimensional manifold $M$ and a Riemannian metric $g$ on $M$ such that $g,g^{-1}\in L^{\infty}$, $g\in W^{1,p}_{\rm loc}(M)$ and such that
@@
    There exists a function $f\in C^{1}(M)$ for which the minimal weak upper gradient and the metric slope (with respect to $\mathsf{d}_{g}$) are strictly larger than the norm of the gradient on a set of positive measure, i.e. $|Df|,|Df|_{w}>|\nabla_{g}f|_{g}$ on a set of positive $\mathrm{vol}_{g}$-measure.
@@
For $d\geq 3$, $p\in[1,d-1)$ there exists a $d$-dimensional manifold $(M,g)$ with $g,g^{-1}\in L^{\infty}_{\rm loc}$ and $g\in W^{1,p}_{\rm loc}(M)$ such that the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is not infinitesimally Hilbertian.
@@
Given a manifold $M$, a Riemannian metric $g$, and a locally Lipschitz continuous function $f:M\to\mathbb{R}$, $|Df|$ will always denote its slope (with respect to the distance $\mathsf{d}_{g}$), $|Df|_{w}$ its minimal weak upper gradient, $\nabla_{g}f$ its gradient with respect to $g$ and $\nabla_{euc}f$ its differential with respect to chosen local coordinates (which will be specified).
@@
When working in $\mathbb{R}^{d}$, $e_{i}$ denotes the $i$-th unit vector, $\mathcal{L}^{d}$ the $d$-dimensional Lebesgue measure, and $Id$ denotes the identity matrix.
For any measure space $(X,\mathfrak{m})$ and $\mathfrak{m}$-measurable set $A\subset X$, $\mathbbm{1}_{A}$ denotes the characteristic function of $A$.
@@ 2.1 First order calculus on metric measure spaces
if $\{x\}$ is not isolated, and 00 otherwise.
We endow $(X,\mathsf{d})$ with a non-negative $\sigma$-finite Borel measure $\mathfrak{m}$, obtaining the metric measure space $(X,\mathsf{d},\mathfrak{m})$. Throughout the rest of this work, we assume that there exists a bounded Borel Lipschitz map $V:X\to[0,\infty)$ such that (cf. [2, Ch. 4])
@@
The discussion in [2], Prop. 5.9 and Def. 5.11 shows the existence of a weak upper gradient $|Df|_{w}$ such that $|Df|_{w}\leq G$ $\mathfrak{m}$-a.e. for all other weak upper gradients $G$. We will call it the minimal weak upper gradient of $f$.
@@ 2.2 Manifolds with continuous Riemannian metrics
Proposition 2.9 ([28], Cor. 4.27).
For a manifold $M$ equipped with a continuous Riemannian metric $g$, the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is infinitesimally Hilbertian.
@@ 3 The distance on manifolds with non-continuous Riemannian metrics
From now on we will consider a smooth $d$-dimensional manifold $M$ with a Riemannian metric $g$ such that $g,g^{-1}\in L^{\infty}_{\rm loc}$. Then $(M,g)$ falls into the class of LIP-Riemannian manifolds (see for example [37, 9]). These manifolds and their metric structures, have been studied by Norris [29], and De Cecco and Palmieri [9], as well as Saloff-Coste [31] and Sturm [34] with emphasis on potential theory for uniformly elliptic operators. Note that LIP-manifolds shall not be confused with metric tensors whose coefficients are locally Lipschitz continuous.
@@
In the case of a Riemannian manifold with $g,g^{-1}\in L^{\infty}$, Norris defines the distance $\mathsf{d}$ via
$\mathsf{d}(x,y):=\sup\left\{\left[w(x)-w(y)\right],\ w\mathrm{\ is\ Lipschitz\ continuous\ and\ }|\nabla_{g}w|_{g}\leq 1\ \mathrm{a.e.}\right\}.$
Moreover, for $\rho\in C_{c}^{\infty}(B_{1}^{\mathbb{R}^{d}}(0),[0,\infty))$, with $\int_{\mathbb{R}^{d}}\rho\,{\rm d}x=1$, define $\rho_{\varepsilon}:=\frac{1}{\varepsilon}\rho(\frac{x}{\varepsilon})$ and $g_{\varepsilon}:=\rho_{\varepsilon}*g$ (for details on the convolution on manifolds see [18, 16]). The metrics $g_{\varepsilon}$ are continuous and hence the distance $\mathsf{d}_{g_{\varepsilon}}$ is given as in (2.3). Norris then proves that for $\varepsilon\to 0$, $\mathsf{d}_{g_{\varepsilon}}$ converges to a distance $\mathsf{d}_{0}$ that is locally equivalent to the underlying Euclidean metric induced from a coordinate chart; in particular, [29, Theorem 3.6] states that $\mathsf{d}_{0}=\mathsf{d}=:\mathsf{d}_{g}$.
From the properties of $g$, it follows that all points outside a null set $N_{g}\subset M$ are Lebesgue points of $g,g^{-1}$. For any null set $N\subset M$, define the set
@@
If $\gamma\in C([0,1],M)$ satisfies $\mathcal{L}^{1}(\{t\in[0,1]:\gamma(t)\in N\})=0$, we say that $\gamma$ is transversal to $N$ and write $\gamma\perp N$.
The reference measure is, as in the continuous case, given by ${\rm d}\mathrm{vol}_{g}=\sqrt{\det g}{\rm d}\mathcal{L}^{d}$. From now on, we will make a further assumption (see [9, (1.13)]):
Assumption 3.1.
The manifold $(M,g)$ satisfies one of the following:
@@
    The manifold $M$ has finite $\mathrm{vol}_{g}$-volume.
@@
    We can cover $M$ with coordinate charts such that the smallest eigenvalue of $g$ is uniformly bounded below in these charts and $M$ is complete or the interior of a complete manifold (where the completeness is with respect to the distance induced by charts).
Finally, De Cecco and Palmieri proved in [9, Theorem (2.18) and Theorem (6.1)] that under Assumption 3.1, it holds $\mathsf{d}_{g}=\boldsymbol{\delta}$, where the distance $\boldsymbol{\delta}$ is given by
@@
and $L_{g}(\gamma)=\int_{0}^{1}|\dot{\gamma}_{t}|_{g}{\rm d}t$. In [8, Theorem 4.4], it is shown that for each null set $N_{0}\subset M$, $\boldsymbol{\delta}=\boldsymbol{\delta}_{0}:=\sup_{N,\mathrm{vol}_{g}(N)=0}\boldsymbol{\delta}_{N\cup N_{0}}$.
@@
Let $\gamma:[0,1]\to M$ be a $\mathsf{d}_{g}$-geodesic such that $\gamma\perp N_{g}$. Then, for almost every $t\in[0,1]$, it holds
@@
Let $M$ be a smooth manifold and $g$ a Riemannian metric such that $g,g^{-1}\in L^{\infty}_{\rm loc}$. Let $\gamma:[0,1]\to M$ be a Lipschitz continuous curve and $t\in[0,1]$ such that the metric speed $|\dot{\gamma}_{t}|$ exists at $t$ and $\gamma$ is differentiable at $t$. Let $c:[0,1]\to M$ be another curve such that $c_{t}=\gamma_{t}$, $c$ is differentiable at $t$ and $\dot{c}_{t}=\dot{\gamma}_{t}$. Then the metric speed of $|\dot{c}_{t}|$ of $c$ at $t$ exists and $|\dot{c}_{t}|=|\dot{\gamma}_{t}|$.
@@
We work in local coordinates in $\mathbb{R}^{d}$ and write $\gamma_{t}=c_{t}=:p$ and $\dot{c}_{t}=\dot{\gamma}_{t}=v$. Furthermore, we may restrict to a small neighbourhood $U$ around $p$ and assume that there exist $\lambda,\Lambda>0$ such that $\lambda Id\leq g\leq\Lambda Id$. Then, by the definition of the derivative, we get that
@@
$0\displaystyle 0\leq\lim_{h\to 0}\frac{\mathsf{d}_{g}(\gamma_{t+h},c_{t+h})}{|h|}\leq\lim_{h\to 0}\sqrt{\Lambda}\frac{|a_{\gamma}(h)|_{euc}+|a_{c}(h)|_{euc}}{|h|}=$ 0 .

### v3

@@ Abstract
This paper investigates the failure of certain metric measure spaces to be infinitesimally Hilbertian or quasi-Riemannian manifolds, by constructing examples arising from a manifold $M$ endowed with a Riemannian metric $g$ that is possibly discontinuous, with $g,g^{-1}\in L^{\infty}_{\mathrm{loc}}$ and $g\in W^{1,p}_{\mathrm{loc}}$ for $p\leq\dim M-1$. footnotetext: MSC2020: 53C20, 53C23
@@
Acknowledgements. The author gratefully acknowledges Andrea Mondino for his supervision and guidance, and thanks Alessandro Cucinotta and Alexander Lytchak for valuable discussions. The author is supported by a postgraduate scholarship from the Mathematical Institute at the University of Oxford. Further financial support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813, which funded the 2025 Trimester Program “Metric Analysis” in Bonn, is also acknowledged with gratitude.
@@ 1 Introduction
Manifolds with (semi)-Riemannian metrics of low regularity (in particular below $C^{2}$) have been of growing interest and importance in the field of geometric analysis. They may arise from gluing constructions, or in the case of Lorentzian signature (,+,,+)(-$\ldots,+$), serve to understand physically relevant phenomena from general relativity (see for instance [5]).
One approach to generalise Riemannian manifolds is to drop any underlying smoothness assumption and study metric measure spaces. A metric measure space is a triplet $(X,\mathsf{d},\mathfrak{m})$, where $(X,\mathsf{d})$ is a complete and separable metric space and $\mathfrak{m}$ is a $\sigma$-finite Borel measure on $X$ (playing the role of the volume measure). For a similar generalisation of Lorentzian manifolds (see [24, 6]).
@@
In this paper however, we will focus on the generalised first order calculus on metric measure spaces. First approaches were made by Cheeger [7] to generalise Rademacher’s theorem to PI-spaces, Hajłasz [21] to prove generalised Sobolev embedding theorems, and Shanmugalingam [34]. Later, Ambrosio, Gigli and Savaré introduced a Sobolev calculus based on the notion of test plans in [1] that they proved to be equivalent to the approaches in [7] and [34], leading to the definition of the Sobolev space $W^{1,2}(X)$, and a Laplace operator.
In [16], Gigli studied conditions under which the Laplacian is linear and proposed the notion of infinitesimal Hilbertianity. A metric measure space is infinitesimally Hilbertian if the associated Sobolev space $W^{1,2}(X)$ is a Hilbert space, or equivalently if the Cheeger energy (a generalisation of the Dirichlet energy to metric measure spaces) is a quadratic form. Then the Laplacian is automatically linear and the theory on Dirichlet forms provides various tools to better understand the Laplacian and the heat flow.
Examples of infinitesimally Hilbertian metric measure spaces include geodesically complete weighted Riemannian manifolds [28], locally $\mathrm{CAT}(\kappa)$-spaces equipped with any Borel measure [12], and sub-Riemannian manifolds equipped with any Borel measure [25]. Relevant notions to study the Cheeger energy are the metric speed of absolutely continuous curves ([3], [23]) and the slope of Lipschitz continuous functions.
[…]
@@
However, there are several relevant classes of Riemannian metrics that are non-continuous, for instance the Geroch-Traschen class [14], i.e., those Riemannian metrics $g$ such that $g,g^{-1}\in L^{\infty}_{\rm loc}(M)$ and $g\in W^{1,2}_{\rm loc}(M)$. These regularity assumptions are minimal to compute the distributional Ricci curvature on $M$ (see [20], [18]) and have been studied in both Riemannian (see for example [19]) and Lorentzian signature (see [35]). Moreover, in the context of potential theory for uniformly elliptic operators, the even broader class of (Riemannian) metrics $g$ such that $g,g^{-1}\in L^{\infty}_{\rm loc}(M)$ has been studied in [33], [36], [31] and [9].
If the metric tensor is only measurable, more precisely an equivalence class of metric tensors that is defined up to local $\mathcal{L}^{d}$-almost everywhere equality, Norris (cf. [31]) and De Cecco-Palmieri (cf. [9]) defined a distance $\mathsf{d}_{g}$ that avoids the difficulty that piecewise smooth curves $\gamma$ might lie on a null set where $g$ is not well-defined (see Section 3).
@@
Under which regularity assumptions on $g$ does $(M,\mathsf{d}_{g},g)$ satisfy that for any Lipschitz continuous function $f:M\to\mathbb{R}$ its slope equals the norm of its gradient almost everywhere, i.e., $|Df|(x)=\limsup_{y\to x}\frac{|f(x)-f(y)|}{\mathsf{d}_{g}(x,y)}=|\nabla_{g}f|_{g}(x)$ almost everywhere?
@@
Under which regularity assumptions on $g$ does the metric speed of “almost every” absolutely continuous curve equal the norm of its derivative, i.e., $\lim_{h\to 0}\frac{\mathsf{d}_{g}(\gamma_{t},\gamma_{t+h})}{|h|}=|\dot{\gamma}_{t}|_{g}$ for almost every $t\in[0,1]$?
@@
Under which regularity assumptions on $g$ is $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ infinitesimally Hilbertian?
@@
It is already known that assuming $g\in C^{0}$ is sufficient for the second and third question to have positive answers (see [4], [30]). Moreover, it is known that for $g\in C^{0}$ and $f\in C^{1}$, the slope $|Df|$ of $f$ equals the norm of its gradient [30]. In [37], Sturm establishes that there exist cases, where the answer to Questions 1 and 2 is negative if one does not assume the metric $g$ to be continuous. Moreover, in [10], De Cecco and Palmieri present one example of a metric $g$ on $\mathbb{R}^{2}$ with $g,g^{-1}\in L^{\infty}$ ans such that on a positive measure set of points, the tangent is a normed space, with a norm that does not arise from an inner product.
Assuming $W^{1,p}_{\rm loc}$-Sobolev regularity with growing exponent $p$ can be seen as a bridge from $L^{\infty}_{\rm loc}$-metrics to $C^{0}$-metrics. Indeed, for $p<p^{\prime}$, it holds $W^{1,p^{\prime}}_{\rm loc}\subset W^{1,p}_{\rm loc}$ and by the Sobolev embedding theorem, a Riemannian metric $g$ on a manifold $M$ with $g\in W^{1,p}_{\rm loc}$, if $p>\dim M$, is automatically continuous. This suggests that additional Sobolev regularity could be a discriminant for Questions 1 to 3.
In this paper, we discuss a related example to [10] that facilitates the study of infinitesimal Hilbertianity, and that allows a refinement to demonstrate that the above questions still may have negative answers if we additionally assume Sobolev regularity of the metric tensor.
Outline of the paper and main results. In Section 2, we will recall some basic notions on the first order calculus on metric measure spaces and their explicit descriptions in the case of a manifold equipped with a continuous Riemannian metric. In Section 3, we recall the definition of the distance in the case of a non-continuous Riemannian metric by Norris and De Cecco-Palmieri, and briefly study absolutely continuous curves and test plans in that setting. In Section 4, we will construct Riemannian metrics $g$ with $g,g^{-1}\in L^{\infty}_{\rm loc}$ such that $(M,\mathsf{d}_{g},g)$ is not a quasi-Riemannian manifold and such that $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is not infinitesimally Hilbertian. We start with metrics that only satisfy $g,g^{-1}\in L^{\infty}_{\rm loc}$, providing an introductory example of a non-quasi Riemannian manifold in Subsection 4.1 and an introductory example of a non-infinitesimally Hilbertian space in Subsection 4.2. These examples are related to the one in [10]. In Subsection 4.3, we refine the previous examples to a Sobolev metric $g$ such that $g,g^{-1}\in W^{1,p}_{\rm loc}\cap L^{\infty}_{\rm loc}$ for some $p\leq\dim M-1$. Questions 1 and 2 are then addressed in Subsection 4.3.1, where we prove:
@@
For $d\geq 3$, $p\in[1,d-1]$ there exists a $d$-dimensional manifold $M$ and a Riemannian metric $g$ on $M$ such that $g,g^{-1}\in L^{\infty}$, $g\in W^{1,p}_{\rm loc}(M)$ and such that
@@
    There exists a function $f\in C^{1}(M)$ for which the minimal weak upper gradient and the metric slope are strictly larger than the norm of the gradient on a set of positive measure, i.e. $|Df|,|Df|_{w}>|\nabla_{g}f|_{g}$ on a set of positive $\mathrm{vol}_{g}$-measure.
@@
For $d\geq 3$, $p\in[1,d-1]$ there exists a $d$-dimensional manifold $(M,g)$ with $g,g^{-1}\in L^{\infty}_{\rm loc}$ and $g\in W^{1,p}_{\rm loc}(M)$ such that the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is not infinitesimally Hilbertian.
@@
Given a manifold $M$, a Riemannian metric $g$, and a locally Lipschitz continuous function $f:M\to\mathbb{R}$, $|Df|$ will always denote its slope, $|Df|_{w}$ its minimal weak upper gradient, $\nabla_{g}f$ its gradient with respect to $g$ and $\nabla_{euc}f$ its differential with respect to chosen local coordinates (which will be specified).
@@
When working in $\mathbb{R}^{d}$, $e_{i}$ will denote the $i$-th unit vector, $\mathcal{L}^{d}$ the $d$-dim Lebesgue measure, and $\mathrm{Id}$ will denote the identity matrix. For $i\in\{1,\ldots,d\}$ we will denote by $x_{i}:\mathbb{R}^{d}\to\mathbb{R}$ the projection on the $i$-th coordinate.
For any measure space $(X,\mathfrak{m})$ and an $\mathfrak{m}$-measurable set $A\subset X$, $\mathbbm{1}_{A}$ will denote the characteristic function of $A$.
@@ 2.1 First order calculus on metric measure spaces
if $\{x\}$ is not isolated, and 0 otherwise.
We endow $(X,\mathsf{d})$ with a non-negative $\sigma$-finite Borel measure $\mathfrak{m}$, obtaining the metric measure space $(X,\mathsf{d},\mathfrak{m})$. Throughout the rest of this work, we assume that there exists a bounded Borel Lipschitz map $V:X\to[0,\infty)$ such that (cf. [1, Section 4])
@@
In the following two definitions, we say $\mathfrak{m}$-measurable functions to refer to equivalence classes of $\mathfrak{m}$-measurable funcrtions up to $\mathfrak{m}$-a.e. equality.
@@
The discussion in [1, Proposition 5.9 and Definition 5.11], shows the existence of a weak upper gradient $|Df|_{w}$ such that $|Df|_{w}\leq G$ $\mathfrak{m}$-a.e. for all other weak upper gradients $G$. We will call it the minimal weak upper gradient of $f$.
@@ 2.2 Manifolds with continuous Riemannian metrics
Proposition 2.9 ([30], Corollary 4.27).
Let $M$ be a manifold equipped with a continuous Riemannian metric $g$. Then the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is infinitesimally Hilbertian.
@@ 3 The distance on manifolds with non-continuous Riemannian metrics
From now on we will consider a smooth $d$-dimensional manifold $M$ with a Riemannian metric $g$ such that $g,g^{-1}\in L^{\infty}_{\rm loc}(M)$. Let $U\subset M$ be open such that $U$ lies in one coordinate patch and let $\psi:U\to\mathbb{R}^{d}$ be a local trivialisation. Then, $g,g^{-1}\in L^{\infty}_{\rm loc}(M)$ means that for every such $U$, the components of $\psi_{*}g,\psi_{*}g^{-1}$ are equivalence classes of locally $\mathcal{L}^{d}$-measurable functions up to $\mathcal{L}^{d}$-almost everywhere equality in $\psi(U)$ and that for every compact set $K\subset\psi(U)$, there exist constants $\lambda,\Lambda>0$ such that $\lambda\mathrm{Id}\leq\psi_{*}g\leq\Lambda\mathrm{Id}$ $\mathcal{L}^{d}$-almost everywhere in $K$. Note that under these assumptions the measures $\mathcal{L}^{d}$ and $\mathrm{vol}_{g}$ are locally equivalent and hence have the same null sets.
Then $(M,g)$ falls into the class of LIP-Riemannian manifolds (see for example [40, 9]). These manifolds and their metric structures, have been studied by Norris [31], and De Cecco and Palmieri [9], as well as Saloff-Coste [33] and Sturm [36] with emphasis on potential theory for uniformly elliptic operators. Note that LIP-manifolds shall not be confused with metric tensors whose coefficients are locally Lipschitz continuous.
@@
In the case of a Riemannian manifold with $g,g^{-1}\in L^{\infty}_{\rm loc}$, Norris defines the distance $\mathsf{d}$ via
$\mathsf{d}(x,y):=\sup\left\{w(x)-w(y),\ w\in C^{0,1}_{\rm loc}(M)\ \mathrm{and}\ |\nabla_{g}w|_{g}\leq 1\ \mathrm{vol}_{g}\mathrm{-a.e.}\right\}.$
Here, $w\in C^{0,1}_{\rm loc}(M)$ means that the function $w$ is locally Lipschitz continuous with respect to the Euclidean distance on charts.
Moreover, for $\rho\in C_{c}^{\infty}(B_{1}^{\mathbb{R}^{d}}(0),[0,\infty))$, with $\int_{\mathbb{R}^{d}}\rho\,{\rm d}x=1$, define $\rho_{\varepsilon}:=\frac{1}{\varepsilon}\rho(\frac{x}{\varepsilon})$ and $g_{\varepsilon}:=\rho_{\varepsilon}*g$ (for details on the convolution on manifolds see [20, 18]). The metrics $g_{\varepsilon}$ are continuous and hence the distance $\mathsf{d}_{g_{\varepsilon}}$ is given as in (2.3). Norris then proves that for $\varepsilon\to 0$, $\mathsf{d}_{g_{\varepsilon}}$ converges to a distance $\mathsf{d}_{0}$ that is locally equivalent to the underlying Euclidean metric induced from a coordinate chart; in particular, [31, Theorem 3.6] states that $\mathsf{d}_{0}=\mathsf{d}=:\mathsf{d}_{g}$.
From the properties of $g$, it follows that all points outside a $\mathrm{vol}_{g}$-null set $N_{g}\subset M$ are Lebesgue points of $g,g^{-1}$. For any $\mathrm{vol}_{g}$-null set $N\subset M$, define the set
@@
Here $C^{0,1}_{\rm loc}([0,1],M)$ refers to locally Lipschitz continuous curves with respect to Euclidean distance induced by the local charts. If $\gamma\in C([0,1],M)$ satisfies $\mathcal{L}^{1}(\{t\in[0,1]:\gamma(t)\in N\})=0$, we say that $\gamma$ is transversal to $N$ and write $\gamma\perp N$.
From now on, we will make a further assumption, that will trivially hold for all the relevant example manifolds in this paper.
Assumption 3.1 ([9], (1.13)).
The manifold $(M,g)$ satisfies the following:
@@
    The manifold $M$ has infinite $\mathrm{vol}_{g}$-volume.
@@
    There exists a constant $h\geq 0$ such that all coordinate charts in our atlas satisfy that the smallest eigenvalue of $g$ is uniformly bounded below $h$ and $M$ is complete or the interior of a complete manifold (where the completeness is with respect to the distance induced by the charts as in [9, (1.9)]).
Assumption 3.1 is relevant to most results in [9], in particular in establishing that two different definitions of the distance are equal. Indeed, De Cecco and Palmieri proved in [9, Theorem (2.18) and Theorem (6.1)] that under Assumption 3.1, it holds $\mathsf{d}_{g}=\boldsymbol{\delta}$, where the distance $\boldsymbol{\delta}$ is given by
@@
and $L_{g}(\gamma)=\int_{0}^{1}|\dot{\gamma}_{t}|_{g}{\rm d}t$. In [8, Theorem 4.4], it is shown that for each $\mathrm{vol}_{g}$-null set $N_{0}\subset M$, $\boldsymbol{\delta}=\boldsymbol{\delta}_{0}:=\sup_{N,\mathrm{vol}_{g}(N)=0}\boldsymbol{\delta}_{N\cup N_{0}}$.
@@
Let $\gamma\in C^{0,1}_{\rm loc}([0,1]\to M)$ be a curve such that $\gamma\perp N_{g}$. Then, for almost every $t\in[0,1]$, it holds
@@
Let $M$ be a smooth manifold and $g$ a Riemannian metric such that $g,g^{-1}\in L^{\infty}_{\rm loc}$. Let $\gamma:[0,1]\to M$ be a Lipschitz continuous curve and $t\in[0,1]$ such that the metric speed $|\dot{\gamma}_{t}|$ exists at $t$ and $\gamma$ is differentiable at $t$. Let $c:[0,1]\to M$ be another curve such that $_{t}=\gamma_{t}, cc$ is differentiable at $t$ and $\dot{c}_{t}=\dot{\gamma}_{t}$. Then the metric speed $|\dot{c}_{t}|$ of $c$ at $t$ exists and $|\dot{c}_{t}|=|\dot{\gamma}_{t}|$.
@@
We work in local coordinates in $\mathbb{R}^{d}$ and write $\gamma_{t}=c_{t}=:p$ and $\dot{c}_{t}=\dot{\gamma}_{t}=v$. Furthermore, we may restrict to a small neighbourhood $U$ around $p$ and assume that there exist $\lambda,\Lambda>0$ such that $\lambda\mathrm{Id}\leq g\leq\Lambda\mathrm{Id}$. Then, by the definition of the derivative, we get that
@@
$0\leq\lim_{h\to 0}\frac{\mathsf{d}_{g}(\gamma_{t+h},c_{t+h})}{|h|}\leq\lim_{h\to 0}\sqrt{\Lambda}\frac{|a_{\gamma}(h)|_{euc}+|a_{c}(h)|_{euc}}{|h|}=0.$